# How turbulence increases the bubble-particle collision rate

### v1

@@ Abstract
Introduction.—Bubble-particle collisions are central to the flotation process, which is widely used e.g. in the mining industry to separate minerals through their attachment to rising bubbles. This intricate process involves a wide range of colloid science disciplines[1, 2, 3] with significant complexity stemming from the interplay of hydrodynamics and physicochemical interactions. The hydrodynamic bubble-particle interaction remains of paramount importance for optimization, as it is the rate-determining step that also sets the conditions for other interactions.
@@
Most of the work on bubble-particle collisions draws on the large body of work on particle-particle collisions [4], in particular the seminal work of [5] for the tracer limit and the model proposed by [6] for the kinetic gas limit at large particle inertia. Such models predict the collision rate without considering the hydrodynamic interaction between the collision partners, but must take into account effects such as the segregation of bubbles and particles in a turbulent flow[7, 8]. However, the significant size difference between the larger bubble and the mineral particles in a typical flotation process, makes it important to account for the flow distortion around the bubble. This can cause an encountering particle to be deflected, reducing the actual number of collisions. For a bubble rising in quiescent flow, this process is deterministic and has been widely studied [9, 10, 11, 12, 13, 14, 2, 15, 16]. In this case, the effect of the hydrodynamic interaction can be captured in terms of collision efficiency, which relates the rate of actual collisions to the encounter rate. The lack of a suitable concept to account for turbulence in this context is widely acknowledged in the literature [17, 18, 19]. Existing approaches are flawed because they combine turbulent encounter rates with collision efficiencies applicable to pure gravitational settling [20, 21, 22, 23, 24, 25]. Others use a Reynolds-type decomposition, which is problematic given the strongly nonlinear dependence of the problem on the flow velocity. Another conceptual inconsistency is that the encounter rate is based on the bubble-particle relative velocity, whereas the collision efficiency is determined by the bubble slip relative to the fluid. Some of these deficiencies are overcome in the model of Refs. [26, 27], which however remains limited to tracer particles.
@@
The core problem is to determine the collision rate of small inertial particles with a finite-size object in a turbulent flow. This is of general relevance to many other applications beyond flotation, such as the collision of cloud droplets [28, 29], the accretion of planetesimals by the collection of dust particles [30, 31], depth filtration [32, 33], or bacterial degradation of marine snow [34, 35]. However, the scarcity of data due to the difficulties in performing the necessary experiments and interface-resolved simulations, has hindered the progress of theoretical research in this area.
In this Letter, we present a combined numerical and theoretical study of the effects of turbulence on bubble-particle collisions using state-of-the-art direct numerical simulations. We adopt homogeneous and isotropic turbulence (HIT), in which we measure how turbulent fluctuation modifies the bubble-particle collision rate along the bubble-rising path. The turbulent flow is characterized by the Taylor Reynolds number, $Re_{\lambda}\equiv\sqrt{15u^{\prime}/(\nu\epsilon)}$, where $u^{\prime}$ denotes the root-mean-square velocity of the turbulence , $\nu$ is the kinematic viscosity, and $\epsilon$ is the mean kinetic energy dissipation rate. The bubble dynamics are governed by the bubble Reynolds number $Re_{b}=2r_{b}\overline{U_{b}}/\nu$, where $\overline{U_{b}}$ is the average bubble rising velocity and $r_{b}$ denotes the bubble radius, and by the turbulence intensity $T_{i}=u^{\prime}/\overline{U_{b}}$.
@@
Model.—The collision rate $K$ of particles with a bubble is determined by $K=\Gamma n_{p}$, where $n_{p}$ is the particle number density and the proportionality coefficient $\Gamma$ ($m^{3}s^{-1}$) is commonly referred to as the collision kernel and is determined by the flow [4]. In quiescent flow, the behaviour is deterministic and the particle grazing trajectory $\Psi_{c}$ determines a stream tube for which all the particles contained inside collide with the bubble. Far enough upstream of the bubble, the grazing trajectory leads to a critical radius $r_{c}$, as illustrated in Fig. 1(a). Thus, the collision kernel in quiescent flow can be expressed as $\Gamma_{q}=\pi r_{b}^{2}U_{b}E_{c}$. Here $E_{c}=r_{c}^{2}/r_{b}^{2}$ denotes the collision efficiency[15, 16], which depends on $Re_{b}$, the bubble-particle size ratio $r_{p}/r_{b}$ and the Stokes number $St_{p}=\tau_{p}/\tau_{f}$, with $\tau_{p}$ the particle response time and $\tau_{f}=2r_{b}/U_{b}$ the timescale of the bubble-particle interaction.
@@
To provide an understanding of the relevant physical mechanisms, we propose a statistical model. We start from the assumption that the flow field in the vicinity of the bubble is approximately stationary and uniform during the bubble-particle interaction. This implies that the temporal scale $\tau_{f}$ is shorter than the correlation timescale of flow fluctuations ($O(\tau_{\eta})$). Additionally, the flow correlation length scale ($10~{}\eta$) should be comparable or larger than the bubble size. Within this ‘frozen turbulence’ assumption, the instantaneous bubble-particle collision process in turbulence can be related back to that observed in quiescent flow. The equivalent steady flow problem is characterised by the magnitude of the instantaneous bubble slip velocity $U_{b}^{\prime}$ with corresponding values of $Re_{b}^{\prime}=Re_{b}U_{b}^{\prime}/\overline{U_{b}}$ and $St_{p}^{\prime}=St_{p}U_{b}^{\prime}/\overline{U_{b}}$.
The entire bubble-particle collision process in a turbulent flow can then be viewed as a superposition of collision events in quiescent flow with varying parameters $Re_{b}^{\prime}$ and $St_{p}^{\prime}$. Based on this, the collision kernel can be derived as follows [36]:
@@
where the local instantaneous particle number density $n_{p}^{\prime}$ is introduced to account for fluctuations in this quantity upstream of the bubble. Importantly $E_{c}(Re_{b}^{\prime},~{}St_{p}^{\prime})$ denotes the collision efficiency in quiescent flow for varying flow parameters, which is deterministic and can be parameterised. The collision kernel is therefore determined by the probability density function (PDF) $f(U_{b}^{\prime})$ of the bubble slip velocity $U_{b}^{\prime}$, which is the main input to the model.
@@
Figure 1: (a) Sketch of the grazing trajectory (red dashed line) in quiescent flow. (b) Mean flow streamlines around the bubble in HIT (solid) and quiescent flow (dashed lines). (c) Trajectories of colliding particles ($r_{p}/r_{b}=0.05,~{}St_{p}=0.04$) in HIT compared to the corresponding grazing trajectories (red lines) in quiescent flow at the same $Re_{b}=120$.
@@
Interface-resolved simulations.—For the simulations to test this model, we use a code based on the lattice Boltzmann method (LBM) to solve the Navier-Stokes equations[37], which has been employed before to study dynamics of point-like [38, 39, 40] and finite-size particles [41, 42] in turbulence. We assume the bubble to be fully contaminated with a no-slip boundary condition, spherical and non-deformable. The coupling between the contaminated bubble and the flow is resolved by the immersed boundary method[43, 44, 45]. Two sets of simulations are carried out[36]: 1) a simplified configuration where a constant bubble velocity is imposed, and 2) simulations with a freely rising bubble, where the bubble-fluid density ratio is $10^{-3}$. Matching practically relevant conditions[46], we consider a moderately turbulent flow ($Re_{\lambda}=32$ and 64) and a bubble Reynolds number $Re_{b}\sim O(100)$. Additionally, we conduct simulations of bubble-particle collisions in a quiescent flow at $Re_{b}$ ranging from 80 to 210. In the flotation process, the mineral particles are significantly smaller than the turbulent dissipation scale $\eta$. We represent these as one-way coupled point particles[36], considering the drag force and the added mass force. The particle response time $\tau_{p}=r_{p}^{2}/(3\beta\nu)$ includes the density ratio $\beta=3\rho_{f}/(\rho_{f}+2\rho_{p})$, where $\rho_{f}$ and $\rho_{p}$ denote the density of the fluid and the particle, respectively. Note that the ratio of turbulent dissipative time scale $\tau_{\eta}$ to $\tau_{f}$ in our simulations is $\sim O(1)$. The collision rate, $K=N_{c}/T$ and thus the collision kernel $\Gamma=K/n_{p}$, is measured by counting the number of particles ($N_{c}$) that collide with the bubble over a time period $T$.
Figure 2: (a) Normalised $E_{c}$ as a function of $St_{p}$ in quiescent flow for various $Re_{b}$ compared to the fits according to Eq. (2) (dashed lines). Inset: scaled PDF of $St_{p}^{\prime}$ as a function of $St_{p}^{\prime}$ for different $St_{p}$.(b) Dimensionless collision kernel vs. $St_{p}$ at $Re_{b}=120$ for bubble with imposed velocity in HIT (symbols) and model (dashed lines). (c) Turbulent collision kernel relative to that in quiescent flow for the imposed velocity bubble. Error bars represent fluctuations between subsets of the data.

### v2

@@ I Introduction
Bubble-particle collisions are central to the flotation process, which is widely used e.g. in the mining industry to separate minerals through their attachment to rising bubbles. This intricate process involves a wide range of colloid science disciplines [1, 2, 3] with significant complexity stemming from the interplay of hydrodynamics and physicochemical interactions. The hydrodynamic bubble-particle interaction remains of paramount importance for optimization, as it is the rate-determining step that also sets the conditions for other interactions.
@@
Most of the work on bubble-particle collisions draws on the large body of work on particle-particle collisions [4], in particular the seminal work of [5] for the tracer limit and the model proposed by [6] for the kinetic gas limit at large particle inertia. Such models predict the collision rate without considering the hydrodynamic interaction between the collision partners, but must take into account effects such as the segregation of bubbles and particles in a turbulent flow[7]. However, the significant size difference between the larger bubble and the mineral particles in a typical flotation process, makes it important to account for the flow distortion around the bubble. This can cause an encountering particle to be deflected, reducing the actual number of collisions. For a bubble rising in quiescent flow, this process is deterministic and has been widely studied [8, 2, 9, 10]. In this case, the effect of the hydrodynamic interaction can be captured in terms of collision efficiency, which relates the rate of actual collisions to the encounter rate. The lack of a suitable concept to account for turbulence in this context is widely acknowledged in the literature [11, 12, 13]. Existing approaches are flawed because they combine turbulent encounter rates with collision efficiencies applicable to pure gravitational settling [14, 15, 16, 17]. Others use a Reynolds-type decomposition, which is problematic given the strongly nonlinear dependence of the problem on the flow velocity. Another conceptual inconsistency is that the encounter rate is based on the bubble-particle relative velocity, whereas the collision efficiency is determined by the bubble slip relative to the fluid. Some of these deficiencies are overcome [18], which however remains limited to tracer particles.
@@
The core problem is to determine the collision rate of small inertial particles with a finite-size object in a turbulent flow. This is of general relevance to many other applications beyond flotation, such as the collision of cloud droplets [19, 20], the accretion of planetesimals by the collection of dust particles [21, 22], depth filtration [23, 24], or bacterial degradation of marine snow [25]. However, the scarcity of data due to the difficulties in performing the necessary experiments and interface-resolved simulations, has hindered the progress of theoretical research in this area. In this paper, we present a combined numerical and theoretical study of the effects of turbulence on bubble-particle collisions using direct numerical simulations. We adopt homogeneous and isotropic turbulence (HIT), in which we measure how turbulent fluctuation modifies the bubble-particle collision rate along the bubble-rising path.
II Statistical Model
@@
The collision frequency $K$ of particles with a bubble is determined by $K=\Gamma n_{p}$, where $n_{p}$ is the particle number density, and the proportionality coefficient $\Gamma$ ($m^{3}s^{-1}$) is commonly referred to as the collision kernel and is determined by the flow [4]. In quiescent flow, the collisions between a spherical bubble of radius $r_{b}$ rising at constant velocity $U_{b}$ with small inertial solid particles are deterministic. The particle grazing trajectory $\Psi_{c}$ determines a ’collision tube’, encompassing all the particles contained inside collide with the bubble, as illustrated in Fig. 1(a). In this case, the collision frequency is determined by the collision number flux: $K=Q_{c}n_{p}$. The collision number flux $Q_{c}$ can be determined by the cross-section area of radius $r_{c}$ limited by the grazing trajectories upstream far from the bubble: $Q_{c}=\pi r_{c}^{2}U_{b}$. Therefore, the collision kernel in quiescent flow can be expressed as $\Gamma_{q}=Q_{c}=\pi r_{b}^{2}U_{b}E_{c}$, where the collision efficiency $E_{c}=r_{c}^{2}/r_{b}^{2}$ measures the ratio of collided particle number to the total encountering particle number. Notice here $E_{c}$ depends on the bubble Reynolds number $Re_{b}=2r_{b}U_{b}/\nu$, the ratio of particle radius $r_{p}$ and $r_{b}$, and the Stokes number $St_{p}=\tau_{p}/\tau_{f}$, with $\tau_{p}$ the particle response time and $\tau_{f}=2r_{b}/U_{b}$ the timescale of the bubble-particle interaction [26, 9, 10].
@@
To provide an understanding of the relevant physical mechanisms of bubble-particle collision in turbulence, we propose a statistical model. We start from the assumption that the flow field in the vicinity of the bubble is approximately stationary and uniform during the bubble-particle interaction. Conceptually, this implies that the temporal scale $\tau_{f}$ is (much) shorter than the correlation timescale of flow fluctuations ($O(\tau_{\eta})$), but it remains to be tested from the data under what conditions exactly this assumption is valid in practice. Additionally, we consider the flow correlation length scale to be comparable to or larger than the bubble size. Within this ‘frozen turbulence’ assumption, the instantaneous bubble-particle collision process in turbulence can be related back to that observed in quiescent flow. Therefore, the equivalent steady flow problem is characterized by the magnitude of the instantaneous bubble slip velocity $U_{b}^{\prime}$ with corresponding values of $Re_{b}^{\prime}=2r_{b}U_{b}^{\prime}/\nu$ and $St_{p}^{\prime}=\tau_{p}/(2r_{b}/U_{b}^{\prime})$, as shown in Fig. 1(b). The entire bubble-particle collision process in a turbulent flow can then be viewed as a superposition of collision events in quiescent flow with varying parameters $Re_{b}^{\prime}$ and $St_{p}^{\prime}$. In the framework, the collision number ${N_{c}|_{\Delta\tau}}$ during a short period $\Delta\tau$ ($O(\tau_{\eta})$) in turbulent flow, with mean particle number density $n_{p}$, is given by $N_{c}|_{\Delta\tau}=\pi r_{b}^{2}E_{c}(Re_{b}^{\prime},St_{p}^{\prime})U_{b}^{\prime}n^{\prime}_{p}\Delta\tau$, where $n_{p}^{\prime}$ is the instantaneous encountering particle number density, which depend on $Re_{b}^{\prime}$ and $St_{p}^{\prime}$. Importantly, $E_{c}(Re_{b}^{\prime},~{}St_{p}^{\prime})$ denotes the collision efficiency in quiescent flow for varying flow parameters, which is deterministic and can be parameterized. Then, the total expected collision number $N_{c}$ during a sufficient long time period $T=\sum\Delta\tau$ is given by
$N_{c}=\pi r_{b}^{2}T\int E_{c}(Re_{b}^{\prime},~{}St_{p}^{\prime})U_{b}^{\prime}f(U_{b}^{\prime})n_{p}^{\prime}\rm d\it U_{b}^{\prime},$ (1)
Based on this, the collision kernel can be derived as follows:
@@
Note that collision efficiency has been extensively studied, and the general dependence of $E_{c}$ on the bubble Reynolds number and particle Stokes number has been established in the literature [8, 5, 10]. The collision kernel is thus determined by the probability density function (PDF) $f(U_{b}^{\prime})$ of the bubble slip velocity $U_{b}^{\prime}$, which remains an open question. Furthermore, we note that the collision kernel will also be affected by the preferential sampling for particles with large $St_{p}$ as this this alters the incoming particle number density.
III Numerical methods and simulations
We perform interface-resolved direct numerical simulations to test this model. The turbulent flow is governed by the incompressible Navier-Stokes equations (NSE), which read:
$\frac{\partial\textbf{u}}{\partial t}+\textbf{u}\cdot{\nabla}\textbf{u}$ $=$ $-\frac{1}{\rho_{f}}{\nabla}p+\nu\ \nabla^{2}\textbf{u}+\textbf{f}+\textbf{f}_{b},$ (3)
${\nabla}\cdot\textbf{u}$ $=$ $0,$ (4)
where $\mathbf{u}$ is the fluid velocity, $p$ denotes the pressure and the parameters are the kinematic viscosity $\nu$ and the reference liquid density $\rho_{f}$. The vector $\mathbf{f}$ denotes an external random large-scale volume force, which is statistically homogeneous and isotropic, with constant-in-time global energy input [27]. This force is used to generate and maintain the turbulent flow. The bubble-free turbulent flow intensity is characterized by the Reynolds number based on the Taylor microscale, $Re_{\lambda}=\sqrt{15u^{\prime}/(\nu\epsilon)}$, where $u^{\prime}$ denotes the root-mean-square velocity of the turbulence and $\epsilon$ is the mean energy dissipation rate. The vector $\mathbf{f}_{b}$ accounts for the bubble-fluid two-way coupling.
@@ III Numerical methods and simulations
Figure 1: (a) Sketch of the grazing trajectory (red dashed line) in quiescent flow. The shaded region indicates the collision tube where all particles collide on the bubble. (b) Sketch of the bubble-particle collision model under temporary bubble slip velocity $U_{b}^{\prime}$. (c) Mean flow streamlines around the bubble for the case of imposed velocity bubble (solid) and quiescent flow (dashed lines) at $\overline{Re_{b}}=120$. (d) Trajectories of colliding particles ($r_{p}/r_{b}=0.05,~{}St_{p}=0.04$) for the case of imposed velocity bubble compared to the corresponding grazing trajectories (red lines) in quiescent flow at the same $Re_{b}=120$. (e) Snapshot of the bubble-particle collision process in turbulence for the imposed-velocity bubble with flow from left to right in the bubble frame of reference. Incoming particles that end up colliding with the bubble are marked in red.
In the practice of flotation, commonly a large amount of surfactants is present in the liquid. Therefore, it is reasonable to assume that the bubbles with a typical Reynolds number below 200 are fully contaminated and approximately spherical, leading to a nearly no-slip boundary condition [2, 10]. The Weber number $We=2r_{b}\rho_{f}U_{b}^{2}/\chi$ based on the surface tension $\chi$ of water and the bubble rise velocity is O(0.1) for our simulations and even lower based on turbulent fluctuations. Therefore bubble deformations remain negligible even if surface tension is lowered by the presence of surfactants. Under these conditions, the bubble behaves similarly to buoyant spheres. In this case, the translation and rotation of the bubble are governed by the Newton-Euler equations (NEE)
$m_{b}\frac{\rm{d}\mathbf{U}_{b}}{{\rm d}t}$ $=$ $\oint_{S_{b}}\bm{\sigma}\cdot\mathbf{n}\ {\rm d}S+V_{b}(\rho_{b}-\rho_{f})\mathbf{g},$ (5)
$\frac{{\rm d}\bm{\mathbf{I}}\mathbf{\Omega}}{{\rm d}t}$ $=$ $\oint_{S_{b}}(\mathbf{x}-\mathbf{x}_{b})\times(\bm{\sigma}\cdot\textbf{n})\ {\rm d}S,$ (6)
where $\mathbf{U}_{b}(t)={\rm d}\mathbf{x}_{b}/{\rm d}t$ and $\mathbf{\Omega}(t)$ respectively are the bubble velocity and angular velocity vectors at position $\mathbf{x}_{b}(t)$. The bubble mass is given by $m_{b}=\rho_{b}V_{b}$ ($\rho_{b}$ the bubble density and $V_{b}$ the volume) and $\bm{\mathbf{I}}$ is the moment of inertia tensor. $\bm{\sigma}=-p\bm{I}+\rho\nu(\nabla\textbf{u}+\nabla\textbf{u}^{T})$ is the fluid stress tensor, $\textbf{x}-\textbf{x}_{b}$ the position vector relative to the bubble center and n the outward-pointing normal to the bubble surface $S_{b}$. This term $\bm{\sigma}$ is solved by the immersed boundary method (IBM) to account for the two-way coupling between the flow and the bubble. In the last decades, IBM has been widely used for studying the multi-phase flow [28, 29, 30]. In IBM, the Euler grids of the fluid are fixed and the surface of the bubble is represented by Lagrangian nodes which move with the bubble motion. To avoid the formation of a mean flow due to the buoyancy driving exerted by the bubble,
we compensate the average force applied from the bubble to the liquid to attain a statistically stationary state [31, 32]. To achieve this, the spatial average of the IBM volume force term needs to be subtracted. More explicitly, we compute the average at each simulation time step:
$\langle\mathbf{f}^{ibm}\rangle_{\Omega}(t)=\frac{1}{||\Omega||}\int_{\Omega}\mathbf{f}^{ibm}(\mathbf{x},~{}t)\rm{d}\mathbf{x},$ (7)
where $\Omega$ denotes the entire computational domain, $\langle\cdots\rangle_{\Omega}$ indicates the spatial average over the entire $\Omega$ region and $||\Omega||$ is the volume of $\Omega$ region. Then, the bubble-related contribution to the volume force, $\mathbf{f}_{b}$, is obtained by
$\mathbf{f}_{b}(\mathbf{x},~{}t)=\mathbf{f}^{ibm}(\mathbf{x}~{},t)-\langle\mathbf{f}^{ibm}\rangle_{\Omega}(t),$ (8)
We adopt an implicit method [33] to solve the bubble dynamics as the conventional explicit method is numerically unstable when the bubble is light [30]. To avoid the built-up of momentum in the triply periodic domain over time, the force applied by the bubble on the liquid is compensated to attain a statistically stationary state [32].
Simulation cases $Re_{\lambda}$ $\eta/\Delta x$ $\tau_{\eta}/\Delta t$ $L/\eta$ $T_{L}/\tau_{\eta}$ $\lambda/\eta$ $\overline{Re_{b}}$ $T_{i}$
imposed-velocity bubble 32 8 780 27.3 9.1 11.6 120 0.25
freely rising bubble 32 8 780 27.3 9.1 11.6 110 0.27
freely rising bubble 64 4 780 67.9 16.7 16.2 150 0.53
Table 1: Parameter of the numerical simulations and relevant turbulence scales. $Re_{\lambda}$: Taylor-Reynolds number, $\eta=(\nu^{3}/\epsilon)^{1/4}$: Kolmogorov dissipation length scale in grid space units $\Delta x$, $\tau_{\eta}$: Kolmogorov time scale in time-step units $\Delta t$, $L=u^{\prime 3}/\epsilon$: integral scale, $T_{L}=L/u^{\prime}$: large-eddy turnover time, $\lambda=u^{\prime}\sqrt{15\nu/\epsilon}$: Taylor micro-scale, $Re_{\lambda}$: Taylor-Reynolds number, $\overline{Re_{b}}=2r_{b}\overline{U_{b}}/\nu$: bubble Reynolds number based on the mean rising velocity $\overline{U_{b}}$, $T_{i}=u^{\prime}/\overline{U_{b}}$: turbulent intensity
@@
We use a code based on the lattice Boltzmann method (LBM) to solve the Navier-Stokes equations [34, 35]. Two sets of simulations are carried out: 1) a simplified ideal configuration where a constant bubble velocity is imposed, and 2) simulations with a freely rising bubble, where the bubble-fluid density ratio is $\rho_{b}/\rho_{f}=10^{-3}$. Matching practically relevant conditions [36], we consider a moderately turbulent flow ($Re_{\lambda}=32$ and 64) and a bubble Reynolds number $Re_{b}\sim O(100)$. Detailed parameters of the simulations are summarized in table 1. There are two computational conditions that should be satisfied in the simulations. Firstly, the wake of the fully contaminated bubble presents a steady axis-symmetric vortex at such a high $Re_{b}$ in a quiescent flow [9, 37]. As the flow domain is periodic, the incoming flow ahead the bubble might be disturbed by the remnants of this bubble wake. To avoid this issue, the flow domain in the bubble rising direction should be sufficiently large. The whole flow domain adopted in this work is rectangular with uniform grid sizes, $N_{x}\times N_{y}\times N_{z}=3072\times 512\times 512$, where the bubble rises in the $x$-direction. We found that using this domain length was sufficient to avoid wake effect. Secondly, the number of lattices within the boundary layer of the bubble should be sufficient. The approximate boundary layer thickness is estimated as $\delta=1.13/Re_{b}^{1/2}$ [37], where $\delta$ denotes the boundary layer thickness normalized by $2r_{b}$. Additionally, the minimum particle size ($r_{p}/r_{b}=0.025$) should be larger than the lattice unit to ensure the accuracy of interpolation when the particle is close to the bubble. To this end, the bubble diameter is resolved by 80 grids to make sure the boundary layer is well resolved, as well as the minimum distance of the point-like particle to the bubble is larger than one grid unit. The ratio of turbulent dissipation time scale $\tau_{\eta}$ to $\tau_{f}$ in our simulations is close to 1.2. We note that the typical auto-correlation time scale of the turbulent flow is longer than $\tau_{\eta}$, which implies the validity of the assumption of frozen turbulence in our model. In the flotation process, the mineral particles are significantly smaller than the turbulent dissipation scale $\eta$. We represent these as one-way coupled point particles, considering the non-Stokesian drag force and the added mass force. We do not account for potential alterations to the drag force when a particle is close to the bubble surface. The particle response time $\tau_{p}=r_{p}^{2}/(3\beta\nu)$ includes the density coefficient $\beta=3\rho_{f}/(\rho_{f}+2\rho_{p})$, where $\rho_{p}$ denote the density of the particle. The particle dynamics is driven by the instantaneous turbulent flow. The collision frequency, $K=N_{c}/T$ and thus the collision kernel $\Gamma=K/n_{p}$, is measured by counting the number of particles ($N_{c}$) that collide with the bubble over a long time period $T$. A collision is detected when the distance between the particle and bubble is smaller than $r_{p}+r_{b}$. Additionally, we conduct simulations of bubble-particle collisions in a quiescent flow at $Re_{b}$ ranging from 80 to 210 to obtain the dependence of $E_{c}$ on $Re_{b}$ and $St_{p}$. In these simulations, a constant-velocity inflow is applied at the inlet, while a homogeneous Neumann boundary condition is imposed at the outlet. The bubble is fixed at the center of the domain and the spatial resolution is kept the same as for the turbulent flow case and the turbulence forcing term $\mathbf{f}$ is switched off.
IV Results